HETAV PATEL

Grade 11,

TANGENTS AND NORMALS find the points on the curve xy + 4 = 0 at which the tangents are inclined at an angle of 45 degree with the x axis.

TANGENTS AND NORMALS 
 find the  points on the curve  xy + 4 = 0 at which the tangents are inclined at an angle of 45 degree with the x axis.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Slope = m
m = tan45 = 1
xy + 4 = 0
y + xy' = 0
y + x = 0
y = -x
Put in curve
x(-x)+4 = 0
x^2 = 4
x = \pm 2
y = \mp 2